OpenStudy (anonymous):

1 ----------dx list steps and notation used please! 1+4x^2

OpenStudy (anonymous):

the integral of that

OpenStudy (anonymous):

yes sorry, didn't know I can draw integrals here

OpenStudy (perl):

$$ \large{ \int \frac{1}{1 + 4x^2}~dx = \int \frac{1}{1 + (2x)^2}~dx } $$

OpenStudy (perl):

you can do u-substitution u = 2x

OpenStudy (anonymous):

\[0.5\arctan2X+C\]

OpenStudy (anonymous):

that's the answer

OpenStudy (perl):

yes

OpenStudy (anonymous):

should I do 1+U^2?

OpenStudy (anonymous):

I didn't understand how to solve it....

OpenStudy (anonymous):

what? how can you multiply the denominator by 2?

OpenStudy (anonymous):

I understand, but why would you do so?

OpenStudy (perl):

$$ \large{ u = 2x\\ du = 2dx\\ \frac{du}{2} = dx\\ ~~~~~~~~~~~\\ \int \frac{1}{1 + (2x)^2}~dx =\int \frac{1}{1 + u^2}\cdot\frac{du}{2} = \frac{1}{2}\cdot \int \frac{1}{1 + u^2} \\ = \frac{1}{2} \arctan(u) + C = \frac{1}{2} \arctan(2x) + C } $$

OpenStudy (anonymous):

Thanks allot! I think I've actually got it!